OpenStudy (anonymous):

Kelly went to visit her uncle's farm. She noticed that the farm has a large number of cows and chickens. She counted 268 legs and 119 heads. How many cows and how many chickens are on Kelly's uncle's farm? Cows: Chickens:

OpenStudy (johnweldon1993):

She counted 268 legs...and 119 heads cows have 4 legs...chickens have 2 each only has 1 head so....you have a system of equations let 'c' = chicken and 'w' = cow 2c + 4w = 268 c + w = 119 can you solve this?

OpenStudy (anonymous):

It cannot be solved.

OpenStudy (anonymous):

And how can I do that?

OpenStudy (johnweldon1993):

2c + 4w = 268 c + w = 119 lets solve that second equation....for the variable 'w'

OpenStudy (johnweldon1993):

c + w = 119 to solve for 'w' we subtract 'c' from both sides w = 119 - c

OpenStudy (anonymous):

could I do c + w = 119 -w -w c = 119-w it's confusing.

OpenStudy (johnweldon1993):

Well yes okay...lets do it your way :) that is how you would solve for 'c'

OpenStudy (anonymous):

but then what could I do? there's still "w" and if I were to divide c by 119 would that give me anything over c/119?

OpenStudy (johnweldon1993):

you have found that c = 119 - w right? Now go back to the FIRST equation... 2c + 4w = 268 plug in what you found for 'c'

OpenStudy (johnweldon1993):

*We will solve for 'w' very soon :) *next step actually! :)

OpenStudy (anonymous):

so... 2 (119-w) +4w =226?

OpenStudy (anonymous):

238-2w +4w=226 -238 -238 -2w +4w = -212 2w= -212 w= -106?

OpenStudy (johnweldon1993):

Actually this is a weird one...but one thing you did wrong above was what is 226 - 238...this = -12 correct? so theoretically we would have 2w = -12 w = -6 Let me do this myself real quick okay?

OpenStudy (johnweldon1993):

ahh nvm....just a simple typo above...where you said "so... 2 (119-w) +4w =226?" 226 should be 268 :)

OpenStudy (anonymous):

lol ohhh ok xD

OpenStudy (johnweldon1993):

haha I was like...THIS IS IMPOSSIBLE!! haha okay so go on from 2 (119-w) +4w =268

OpenStudy (anonymous):

so it would be 238-2w +4w= 268 -238 -238 2w=30 w=15

OpenStudy (johnweldon1993):

there we go! lol so we have 15 cows.....now...we solve for 'c' remember c + w = 119 we now know w = 15 so... c + 15 = 119 what is 'c' ?

OpenStudy (anonymous):

wouldn't c equal 104? :D

OpenStudy (johnweldon1993):

it would indeed! now check to make sure it is correct take the second equation 2c + 4w = 268 plug in your values for 'c' and 'w' to make sure it comes out correct! :)

OpenStudy (anonymous):

2(104)+ 4(15)=268 208+60 omggg <333

OpenStudy (johnweldon1993):

haha there you go! :D see, easy! :)

OpenStudy (anonymous):

lol! thank you so much (: <3

OpenStudy (johnweldon1993):

anytime! :)